OpenStudy (anonymous):

Determine the domain of the function f(x)=√1-x All real numbers x > 1 x ≤ 1 All real numbers except 1

ganeshie8 (ganeshie8):

the thing inside radical must always be \(\ge\) 0

OpenStudy (jdoe0001):

\(\bf \sqrt{1-x} \ \ ?\)

ganeshie8 (ganeshie8):

1-x \(\ge\) 0

ganeshie8 (ganeshie8):

solve x

OpenStudy (anonymous):

ok thank you

ganeshie8 (ganeshie8):

np:) wat do u get after solving x ?

OpenStudy (anonymous):

im solving it right now

ganeshie8 (ganeshie8):

1-x \(\ge\) 0 -1 -1

ganeshie8 (ganeshie8):

-x \(\ge\) -1 x \(\le\) 1

ganeshie8 (ganeshie8):

so the domain is x \(\le\) 1

OpenStudy (anonymous):

yep thats what I GOT:)

ganeshie8 (ganeshie8):

good :)